# references.bib

### v1

@@
Solid state quantum emitters are a prime candidate in distributed quantum technologies since they inherently provide a spin-photon interface. An ongoing challenge in the field, however, is the low photon extraction due to the high refractive index of typical host materials. This challenge can be overcome using photonic structures. Here, we report the integration of V2 centers in a cavity-based optical antenna. The structure consists of a silver-coated, -thin 4H-SiC membrane functioning as a planar cavity with a broadband resonance yielding a theoretical photon collection enhancement factor of 34. The planar geometry allows us to identify over 20 single V2 centers at room temperature with a mean (maximum) count rate enhancement factor of 9 (15). Moreover, we observe 10 V2 centers with a mean absorption linewidth below at cryogenic temperatures. These results demonstrate a photon collection enhancement that is robust to the lateral emitter position.
@@ INTRODUCTION
Optically active spins in solid-state materials are promising candidates for applications in quantum technologies [Awschalom2013, Awschalom2018, Atatuere2018]. Besides the well-established diamond platform [Jelezko2006, Doherty2013], 4H-silicon carbide (SiC) has gained growing attention in the field. It hosts several promising emitters such as the silicon-vacancy center (V) [Kraus2014, Widmann2014, Nagy2019], the divacancy center (VV) [Son2006, Christle2017, Li2022], as well as the more recently investigated nitrogen-vacancy center [VonBardeleben2015, Wang2020] and the vanadium center [Wolfowicz2020, Cilibrizzi2023]. This has yielded a number of encouraging results with regards to quantum networks, such as single shot readout of the electron spin state, through charge state conversion using the VV center [Anderson2022], or through a nearby nuclear spin in case of the V center [Lai2024, Hesselmeier2024], as well as spin-photon entanglement using the V [Fang2024].
In addition, silicon carbide offers excellent nanofabrication capabilities. This enables implementing various photonic structures that improve the rate of collected photons from emitters through their geometry, e.g., solid-immersion lenses (SIL) [Sardi2020, Bekker2023], nanopillars [Radulaski2017], and waveguides [Babin2021, Krumrein2024]. It also allows for the fabrication of more complex structures such as photonic crystal cavities [Bracher2017, Lukin2020], or disk resonators [Wang2021, Lukin2023] that can further boost the photon rate through Purcell enhancement. While nanophotonic cavities offer the highest enhancement among the above mentioned structures, they rely on complicated fabrication processes, tight restrictions on the emitter placement, and complex methods to keep the emitter on resonance with the structure.
@@
A schematic of the cavity antenna is shown in Figure 1(a). The structure is based on a sub-micron-thin 4H-SiC membrane with silver coatings on both sides. Due to the high reflectance of the silver, the structure acts as a Fabry-Pérot cavity. Choosing the thickness of the upper silver layer well below makes it transparent enough such that the cavity modes can be converted to free-propagating modes in the upper direction. The lower silver layer is kept thick enough (several ) to fully reflect the light. As already discussed in [Fuchs2021], this results in a higher photon emission rate and an increased directivity towards the collectable solid angle of an objective for the electric dipole radiation of a color center inside the SiC membrane. Finally, the upper silver layer is capped by an SiO layer to prevent fast oxidation of the silver at ambient conditions. Its thickness is optimized to work as an anti-reflective coating [Fuchs2021]. Note, that plasmonic resonances at the interfaces are a loss channel for this structure rather than the enhancement mechanism like in other approaches [Shalaginov2019, Zhou2023].
@@
For a given set of antenna layer thicknesses, the wavelength dependent emission enhancement over an emitter in bulk crystal is shown as the orange curve in Figure 1(b). We numerically optimize the thickness of the upper silver layer, the membrane, the SiO capping layer, and the distance of the dipole from the lower silver layer to yield a maximum total enhancement [Software2021] (see Supporting Information for details). The total enhancement is calculated by weighting the wavelength dependent enhancement over the normalized fluorescence spectrum of the V2 center, shown in blue in Figure 1(b). Figure 1(c) shows the total enhancement for the collection within a spectral band of as a function of the two most relevant thicknesses in the design, i.e., the upper silver layer and the membrane thickness. We choose this specific wavelength range, since our detector becomes much less efficient for wavelengths above (see Supporting Information). The structure yields a significant enhancement for membrane thicknesses between and upper silver thicknesses between with a maximum enhancement factor of 34. To study the influence of the dipole position within the thin SiC membrane, we simulate the enhancement factor as a function of the membrane thickness and the relative dipole position at the optimized upper silver layer thickness of in Figure 1(e). For the first cavity mode, the enhancement is very robust to a displacement of the dipole from its optimal position in the middle of the membrane. It stays in between for a relative dipole position in the membrane in between from the lower silver-SiC interface. The robustness of the enhancement on more parameters is shown in the Supporting Information.
@@
The fabrication of the antenna structure starts with the steps described in our previous work [Heiler2024] in order to create sub-µm thin 4H-SiC membranes with integrated V2-centers. With this, we create a sample with a total thickness of containing several subregions with a thickness below , which we refer to as membranes in the following. In the last step of this process, we etch our sample down to a thickness of in the central part of one membrane, measured by a home-built white light interferometer (see Supporting Information for details on the membrane thickness). The membrane is subsequently coated using electron beam evaporation with a target silver thickness of for the thick silver layer (full-reflective mirror) and for the thin silver layer (leaky mirror). In the final step, the SiO layer is sputtered on top of the thin silver layer. To check the resonance of the fabricated structure, we measure its reflectivity spectrum after the coating steps using the WLI. The result, normalised by a reference spectrum from a mirror, is depicted in Figure 1(d) and shows a reflection dip at that covers a big part of the phonon side band of the V2 center. We can model the measured data using a transfer matrix model and find a good agreement for an upper silver layer of and a SiC thickness of , confirming the earlier WLI measurement.
@@
For better statistics on the enhancement factor, we investigate a higher number of emitters. At first, we record ODMR for individual spots on the bulk and antenna sample to find V2 centers. We set a threshold of at least contrast in the ODMR to use an emitter for further measurements, since a much lower contrast compared to the average indicates fluorescence from different defects, e.g., the V1 center, or crystal damage in the spot. With this, we find that of the bright spots in the antenna are V2 centers which is lower compared to in the bulk reference sample. We suspect the close proximity to the surface in our antenna sample yields a higher number of surface related defects in confocal scans, thus, decreases the relative number of V2 emitters. However, with a mean density of we still find a high number of V2 centers in the antenna sample. For a faster identification of V2 centers among the bright spots, we use a pre-selection based on the polarization of the emission (see Supporting Information for details). In the following, we use 25 representative emitters in the antenna, 23 in the bulk and 7 emitters in the SIL sample.
@@
At resonant excitation with low power of (), we find emitters with a stable PLE over a long time scale, as shown for a representative emitter in 3(a) for half an hour. We perform excitation power dependent PLE measurements to estimate the absorption linewidths of 10 arbitrarily selected emitters out of the emitters with distinguishable transitions. Each emitter is measured for 10 single sweeps at up to six different excitation powers in between . The mean A and A linewidths for each power obtained by averaging the linewidth fit parameter from a double peak Gaussian fit of each single sweep are depicted in 3(b). For comparison, the mean A and A linewidths of an emitter in a SIL sample are measured as well, using a spin-mixing microwave signal. As expected due to the much larger distance of the emitter from the SiC surface, the linewidths from the SIL emitter lie below the respective antenna sample transition linewidths, except for the highest excitation power, where power broadening becomes significant for both structures. In total, the average linewidth for each power for the antenna emitters lies below . Data on the power dependent linewidths of each individual antenna emitter can be found in the Supporting Information.
@@
Finally, we investigate the count rate enhancement of the structure under resonant excitation. To prevent a (partial) spin-polarization, which would lower the amount of collected photons, we scan two lasers separated by simultaneously to drive the A and A transition at the same time (details in the Supporting Information). With this excitation scheme, we record power-dependent PLE for each three emitters in the reference bulk and SIL sample and 10 emitters in the antenna sample up to an excitation power where the emitter ionizes already after a single PLE scan most of the time. We evaluate the photon count rate for each excitation power by averaging the amplitudes from Gaussian fits on the single PLE lines. However, we find that especially for the emitters in the antenna sample, the power dependent average count rates do not reach the typical non-linear regime of saturation since they ionize in almost every line before. This effect results in saturation curves where no saturation power can be confidently extracted via a fit apart from 4 antenna emitters (see Supporting Information).
@@
Compared to nanopillars, SILs, or nanobeam resonators that enhance emitters in a small spatial region, our structure offers the big advantage to enhance emitters in a two-dimensional plane, that is limited in size solely by the fabrication. Thus, further work on the fabrication, i.e. creating all color centers in the optimal depth using ion-implantation methods [Ohshima2018, Wang2019] and further optimizing the thickness homogeneity, can be used to significantly scale up the number of enhanced emitters in a single structure. This simplifies the integration of defects with a low creation yield, e.g., SiC divacancies near stacking faults [He2024]. Moreover, if the structure is transferred to a system with emitters that have a broad inhomogeneous linewidths, e.g., optically active molecules [Checcucci2017, Toninelli2021] or rare earth ion quantum emitters [Thiel2011], the big number of enhanced emitters eases finding emitters with matching emission lines, enabling two-emitter entanglement. The versatile design also enables enhancing the ZPL using the spectrally smaller resonance of a thicker membrane and making the structure very appealing in the context of quantum communication where indistinguishable photons are key.
@@
The data supporting the presented findings are available from the corresponding author upon reasonable request.

### v2

@@
Solid state quantum emitters are a prime candidate in distributed quantum technologies since they inherently provide a spin-photon interface. An ongoing challenge in the field, however, is the low photon extraction due to the high refractive index of typical host materials. This challenge can be overcome using photonic structures. Here, we report the integration of V2 centers in a cavity-based optical antenna. The structure consists of a silver-coated, -thin 4H-SiC membrane functioning as a planar cavity with a broadband resonance yielding a theoretical photon collection enhancement factor of . The planar geometry allows us to identify over 20 single V2 centers at room temperature with a mean (maximum) count rate enhancement factor of 9 (15). Moreover, we observe 10 V2 centers with a mean absorption linewidth below at cryogenic temperatures. These results demonstrate a photon collection enhancement that is robust to the lateral emitter position.
@@ INTRODUCTION
Optically active spins in solid-state materials are promising candidates for applications in quantum technologies [Awschalom2013, Awschalom2018, Atatuere2018]. Besides the well-established diamond platform [Jelezko2006, Doherty2013], 4H-silicon carbide (4H-SiC) has gained growing attention in the field. It hosts several promising emitters such as the silicon-vacancy center (V) [Kraus2014, Widmann2014, Nagy2019], the divacancy center (VV) [Son2006, Christle2017, Li2022], as well as the more recently investigated nitrogen-vacancy center [VonBardeleben2015, Wang2020], and the vanadium center [Wolfowicz2020, Cilibrizzi2023]. This has yielded a number of encouraging results with regards to quantum networks, such as single shot readout of the electron spin state, through charge state conversion using the VV center [Anderson2022], or through a nearby nuclear spin in case of the V center [Lai2024, Hesselmeier2024], as well as spin-photon entanglement using the V [Fang2024].
In addition, silicon carbide offers excellent nanofabrication capabilities. This enables implementing various photonic structures that improve the rate of collected photons from emitters through their geometry, e.g. solid-immersion lenses (SIL) [Sardi2020, Bekker2023], nanopillars [Radulaski2017], and waveguides [Babin2021, Krumrein2024]. It also allows for the fabrication of more complex structures, such as photonic crystal cavities [Bracher2017, Lukin2020] or disk resonators, [Wang2021, Lukin2023] that can further boost the photon rate through Purcell enhancement. While nanophotonic cavities offer the highest enhancement among the above mentioned structures, they rely on complicated fabrication processes, tight restrictions on the emitter placement, and complex methods to keep the emitter on resonance with the structure.
@@
A schematic of the cavity antenna is shown in Figure 1(a). The structure is based on a sub-micron-thin 4H-SiC membrane with silver coatings on both sides. Due to the high reflectance of the silver, the structure acts as a Fabry-Pérot cavity. Choosing the thickness of the upper silver layer well below makes it transparent enough such that the cavity modes can be converted to free-propagating modes in the upper direction. The lower silver layer is kept thick enough (several ) to fully reflect the light. This results in a higher photon emission rate and an increased directivity towards the collectable solid angle of an objective for the electric dipole radiation of a color center inside the SiC membrane. Finally, the upper silver layer is capped by an SiO layer to prevent fast oxidation of the silver at ambient conditions. Its thickness is optimized to work as an anti-reflective coating [Fuchs2021]. Note, that plasmonic resonances at the interfaces are a loss channel for this structure rather than the enhancement mechanism like in other approaches [Shalaginov2019, Zhou2023].
@@
For a given set of antenna layer thicknesses, the wavelength dependent emission enhancement over an emitter in bulk crystal is shown as the orange curve in Figure 1(b). We numerically optimize the thickness of the upper silver layer, the membrane, the SiO capping layer, and the distance of the dipole from the lower silver layer to yield a maximum total enhancement [Software2021] (see Supporting Information for details). The total enhancement is calculated by weighting the wavelength dependent enhancement over the normalized fluorescence spectrum of the V2 center, shown in blue in Figure 1(b). Figure 1(c) shows the total enhancement for the collection within a spectral band of as a function of the two most relevant thicknesses in the design, i.e., the upper silver layer and the membrane thickness. We choose this specific wavelength range, since our detector becomes much less efficient for wavelengths above (see Supporting Information). The structure yields a significant enhancement for membrane thicknesses between and upper silver thicknesses between with a maximum enhancement factor of . To study the influence of the dipole position within the thin SiC membrane, we simulate the enhancement factor as a function of the membrane thickness and the relative dipole position at the optimized upper silver layer thickness of in Figure 1(e). For the first cavity mode, the enhancement is very robust to a displacement of the dipole from its optimal position in the middle of the membrane. It stays in between for a relative dipole position in the membrane in between from the lower silver-SiC interface. The robustness of the enhancement on more parameters is shown in the Supporting Information.
@@
The fabrication of the antenna structure starts with the steps described in our previous work [Heiler2024] in order to create sub-µm thin 4H-SiC membranes with integrated V2-centers. With this, we create a sample with a total thickness of containing several subregions with a thickness below , which we refer to as membranes in the following. In the last step of this process, we etch our sample down to a thickness of in the central part of one membrane (see Supporting Information for details on the membrane thickness), measured by a home-built white light interferometer (WLI). The membrane is subsequently coated using electron beam evaporation with a target silver thickness of for the thick silver layer (full-reflective mirror) and for the thin silver layer (leaky mirror). In the final step, the SiO layer is sputtered on top of the thin silver layer. To check the resonance of the fabricated structure, we measure its reflectivity spectrum after the coating steps using the WLI. The result, normalised by a reference spectrum from a mirror, is depicted in Figure 1(d) and shows a reflection dip at that covers a big part of the phonon side band of the V2 center. We can model the measured data using a transfer matrix model and find a good agreement for an upper silver layer of and a SiC thickness of , confirming the earlier WLI measurement.
@@
For better statistics on the enhancement factor, we investigate a higher number of emitters. At first, we record ODMR for individual spots on the bulk and antenna sample to find V2 centers. We set a threshold of at least contrast in the ODMR to use an emitter for further measurements, since a much lower contrast compared to the average indicates fluorescence from different defects, e.g., the V1 center, or crystal damage in the spot. With this, we find that of the bright spots in the antenna are V2 centers which is lower compared to in the bulk reference sample. We suspect the close proximity to the surface in our antenna sample yields a higher number of surface related defects in confocal scans, thus, decreases the relative number of V2 emitters. However, with a mean density of we still find a high number of V2 centers in the antenna sample. For a faster identification of V2 centers among the bright spots, we use a pre-selection based on the polarization of the emission (see Supporting Information for details). In the following, we use 25 representative emitters in the antenna, 23 in the bulk, and 7 emitters in the SIL sample.
@@
At resonant excitation with low power of (), we find emitters with a stable PLE over a long time scale, as shown for a representative emitter in Figure 3(a) for half an hour. We perform excitation power dependent PLE measurements to estimate the absorption linewidths of 10 arbitrarily selected emitters out of the emitters with distinguishable transitions. Each emitter is measured for 10 single sweeps at up to six different excitation powers in between . The mean A and A linewidths for each power obtained by averaging the linewidth fit parameter from a double peak Gaussian fit of each single sweep are depicted in Figure 3(b). For comparison, the mean A and A linewidths of an emitter in a SIL sample are measured as well, using a spin-mixing microwave signal. As expected due to the much larger distance of the emitter from the SiC surface, the linewidths from the SIL emitter lie below the respective antenna sample transition linewidths, except for the highest excitation power, where power broadening becomes significant for both structures. In total, the average linewidth for each power for the antenna emitters lies below . Data on the power dependent linewidths of each individual antenna emitter can be found in the Supporting Information.
@@
Finally, we investigate the count rate enhancement of the structure under resonant excitation. To prevent a (partial) spin-polarization, which would lower the amount of collected photons, we scan two lasers separated by simultaneously to drive the A and A transition (details in the Supporting Information). With this excitation scheme, we record power-dependent PLE for three emitters in the reference bulk, three emitters in the SIL, and 10 emitters in the antenna sample up to an excitation power where the emitter ionizes already after a single PLE scan most of the time. We evaluate the photon count rate for each excitation power by averaging the amplitudes from Gaussian fits on the single PLE lines. However, we find that especially for the emitters in the antenna sample, the power dependent average count rates do not reach the typical non-linear regime of saturation since they ionize in almost every line before. This effect results in saturation curves where no saturation power can be confidently extracted via a fit apart from 4 antenna emitters (see Supporting Information).
@@
Compared to nanopillars, SILs, or nanobeam resonators that enhance emitters in a small spatial region, our structure offers the big advantage to enhance emitters in a two-dimensional plane, that is limited in size solely by the fabrication. Thus, further work on the fabrication, i.e. creating all color centers in the optimal depth using ion-implantation methods [Ohshima2018, Wang2019] and further optimizing the thickness homogeneity, can be used to significantly scale up the number of enhanced emitters in a single structure. This simplifies the integration of defects with a low creation yield, e.g., SiC divacancies near stacking faults [He2024]. Moreover, if the structure is transferred to a system with emitters that have a broad inhomogeneous linewidth, e.g., optically active molecules [Checcucci2017, Toninelli2021] or rare earth ion quantum emitters [Thiel2011], the big number of enhanced emitters eases finding emitters with matching emission lines, enabling two-emitter entanglement. The versatile design of the cavity antenna also enables enhancing the ZPL using the spectrally smaller resonance of a thicker membrane and making the structure very appealing in the context of quantum communication where indistinguishable photons are key.
@@
The data supporting the presented findings are available at https://doi.org/10.18419/darus-4262.